Complete the ordered pair so that each is a solution for the given equation.
3x + 4y = 12 (0, ), ( ,3/4), (,0), (8/3, )

Answer by rmromero(383)


To complete the ordered pairs, we need to substitute the given coordinate to the equation.
3x + 4y = 12, (x, y)


When x = 0
3(0) + 4y = 12
4y = 12
y = 3 -------->> (0, 3)


When y = 3/4
3x + 4(3/4) = 12
3x + 3 = 12
3x = 9
x = 3 ----->> (3, (3/4))


When y = 0
3x + 4(0) = 12
3x = 12
x = 4 ---------> (4, 0)


When x = 8/3
3(8/3) + 4y = 12
8 + 4y = 12
4y = 4
y = 1 -------->> ((8/3), 1)